OpenStudy (anonymous):

a sample of a substance that has a density of .824 g/ml has a mass of .451 g. calculate the volume of the sample

OpenStudy (anonymous):

d=m/v 0.824 = 0.451/v Find v.

OpenStudy (anonymous):

that would be 1.827 right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

what wouldnt it be .547 because V=m/density?

OpenStudy (anonymous):

No. d=m/v so v=d/m

OpenStudy (anonymous):

no because if i put 1.827 in the left side of the equation it wont equal .824

OpenStudy (anonymous):

oh wait nevermind, im getting myself confused. lol thanks

OpenStudy (anonymous):

Oh my bad, the answer is 0.451/0.824, but the formula is correct.

OpenStudy (anonymous):

So yes, it is 0.547.

OpenStudy (anonymous):

I mixed it up when entering it into my calculator.

OpenStudy (anonymous):

no it is 1.827 becuase d =.824 and m =.451 so .824/.451=1.827

OpenStudy (anonymous):

no wait, im confused

OpenStudy (anonymous):

OKAY.  so it is 0.451/0.824, so the answer is 0.547.